# Achieving Near-Optimal Regret for Bandit Algorithms with Uniform Last-Iterate Guarantee

### v1

@@ 2.3 New Metric: Uniform Last-Iterate Guarantee
To address the aforementioned issue, we introduce a new performance measure, called uniform last-iterate (ULI), formally defined below.
@@
[…]
According to Definition 2.5, the instantaneous suboptimality of any algorithm with the ULI guarantee can be bounded by a function $F_{\text{ULI}}(\delta,t)$. Moreover, $F_{\text{ULI}}(\delta,t)$ decreases monotonically for large $t$ if its power on $1/t$ is strictly positive, which captures the convergence rate of $\Delta_{A_{t}}$.
Note that the convergence rate of $\Delta_{A_{t}}$ is mainly determined by the power on $1/t$ in $F_{\text{ULI}}$. Moreover, as we will show shortly, an algorithm with the ULI guarantee automatically has a small regret bound. We therefore provide the following lower bound on $F_{\text{ULI}}$.
@@
[…]
@@
[…]
@@
According to Theorem 2.7, if an algorithm is with near-optimal ULI guarantee (i.e., $\kappa=\frac{1}{2}$), then it implies the near-optimality for uniform-PAC bound (the first bullet) and regret bound (the second bullet). Note that it is not the other way around, in the sense that an algorithm with a near-optimal uniform-PAC bound does not necessarily enjoy a near-optimal ULI guarantee as shown in Section 3.2. The proof of Theorem 2.7 can be found in Appendix B.
3 Achieving ULI in Finite Arm Space
In this section, we answer the question whether ULI is achievable. To this end, we examine three common types of bandit algorithms, including elimination-based algorithms, optimistic algorithms, and high-probability adversarial algorithms, in the finite arm setting in which $|{\mathcal{A}}|=K$.
@@ 3.1 Elimination Framework Achieving Near-optimal ULI Guarantee
To examine whether the elimination-type algorithms can achieve the ULI guarantee, we first provide an elimination framework (see Algorithm 1) that ensures the ULI guarantee, and we then show that most elimination-based algorithms fall into this framework. The following result shows that with a proper function $f$ and a positive constant $\beta$, such an elimination framework ensures the ULI guarantee.
@@
[…]
@@
Theorem 3.1 suggests that Eq. (1) is a sufficient condition for elimination-based algorithms to achieve the ULI guarantee. Now, we show that existing elimination algorithms indeed fall into such a framework. We here consider successive elimination (SE) and phased elimination (PE). Notice that we consider SE only for the MAB setting (called SE-MAB, e.g., Algorithm 3 in (Even-Dar et al., 2006)) but PE for both the MAB setting (called PE-MAB, e.g., exercise 6.8 in (Lattimore and Szepesvári, 2020)) and the linear bandit setting (called PE-linear, e.g., Algorithm 12 of Chapter 22 in (Lattimore and Szepesvári, 2020)). Since those algorithms are standard, we defer their pseudocodes to Appendix D. Given Theorem 3.1, the following results show that the elimination framework can be instantiated by these algorithms with proper functions and therefore they achieve the ULI.
@@
For any fixed $\delta\in(0,1)$, elimination framework in Algorithm 1 can be instantiated by the algorithm
  • SE-MAB for MAB to achieve the ULI with
[…]
  • PE-MAB for MAB to achieve the ULI with
[…]
  • PE-linear for linear bandit setting with $K$ arms to achieve the ULI with
[…]
@@ 3.2 Lower Bound for Optimistic Algorithms
In this subsection, we consider the optimistic algorithms and present a lower bound to show that this type of algorithms yield a worse instantaneous performance, compared with elimination-based algorithms. The procedure of optimistic algorithms is summarized as follows. After playing each arm once, at each round $t$, the algorithm plays an arm $A_{t}$ that satisfies
@@
where $N_{a}(t)$ is the number of plays of arm $a$ before round $t$, $\widehat{\mu}_{a}(N_{a}(t))=\frac{1}{N_{a}(t)}\sum_{s=1}^{t-1}X_{s,A_{s}}\mathbb{I}{\left\{{A_{s}=a}\right\}}$ is the empirical mean of arm $a$ after $N_{a}(t)$ times play, and $U_{\delta}(N_{a}(t))$ is a positive bonus function which encourages the exploration.
We first consider optimistic algorithms e.g., upper confidence bound (UCB) (Lattimore and Szepesvári, 2020, Algorithm 3, Chapter 7), which enjoy (near)-optimal regret bounds. This type of algorithms typically uses the bonus function in the form of $\sqrt{\log(2t^{2}/\delta)/N_{a}(t)}$444 We slightly adjust the bonus function to ensure that with probability at least $1-\delta$, the confidence bound holds for all $t\in\mathbb{N}$.. However, the $\log t$ term forces the algorithm to play suboptimal arms infinitely often, and thus they cannot achieve the ULI guarantee. Similarly, other variants (Audibert and Bubeck, 2010; Degenne and Perchet, 2016) with $\log t$ in bonus function, also cannot achieve the ULI guarantee.
We then consider another optimistic-type algorithm, lil’UCB (Jamieson et al., 2014) which obtains the order-optimal sample complexity and avoids $\log(t)$ term in bonus function $U_{\delta}(N_{a}(t))$. The bonus function of lil’UCB is as $\sqrt{\log\left(\delta^{-1}\log_{+}\left(N_{a}(t)\right)\right)/N_{a}(t)}$ where $\log_{+}(x)=\log\left(\max\left\{x,e\right\}\right)$. This bonus function guarantees the number of plays of suboptimal arms is finite. Our main result for lil’UCB is presented as follows. The full description and theoretical analysis of lil’UCB are deferred to Appendix E.
@@
There exists a constant $\alpha\in(0,1)$ that for any $\Delta\in(0,\alpha)$, running lil’UCB on the two-armed bandit instance with deterministic rewards and arm gap $\Delta$ gives $\exists t=\Omega\left(\Delta^{-2}\right)$ such that
$\Delta_{A_{t}}=\Omega\left(t^{-\frac{1}{4}}\sqrt{\log\log\left(\Delta^{-2}\log(\delta^{-1})\right)+\log(\delta^{-1})}\right).$
Theorem 3.3 shows that lil’UCB is not near-optimal ULI, but it is unclear whether it can achieve the ULI guarantee. Recall from Theorem 3.2 that the elimination-based algorithms ensure that with high probability, $\forall t$, $\Delta_{A_{t}}=\widetilde{\mathcal{O}}\left(t^{-\nicefrac{{1}}{{2}}}\right)$. Theorem 3.3 suggests that the convergence rate of lil’UCB is strictly worse than that of elimination-based algorithms, even if it enjoys a near-optimal cumulative performance (Jamieson et al., 2014). In fact, this lower bound holds for all optimistic algorithms as long as the bonus function is in a similar form.
3.3 Meta-algorithm Enabling Adversarial Algorithms to Achieve ULI
Algorithm 2 Meta-algorithm towards ULI
Input: $\delta\in(0,1)$, finite arm set ${\mathcal{A}}$, and base-algorithm $\mathtt{Alg}$ with function $g(\delta)$ that satisfies Condition. 3.4.
Initialize: ${\mathcal{A}}_{1}=[K]$.
for $m=1,2,\ldots$ do
In this subsection, we propose a meta-algorithm shown in Algorithm 2 (see Appendix F) which enables any high-probability adversarial algorithm, with a mild condition, to achieve the ULI guarantee. Then, we show that existing high-probability adversarial bandit algorithms naturally meet this condition in both MAB and linear bandit settings. For notational simplicity, we follow the convention of adversarial analysis to use loss $\ell_{t,a}=1-X_{t,a}$. Note that all algorithms in this subsection require $[0,1]$-boundedness assumption on loss.
Condition 3.4.
In fact, many high-probability adversarial bandit algorithms naturally meet Condition. 3.4, but require stronger analysis. In particular, we show that Exp3.P (Auer et al., 2002b) for MAB holds this condition with $g(\delta)=\mathcal{O}\big{(}|{\mathcal{A}}|\log(|{\mathcal{A}}|/\delta)\big{)}$ (omitted proof can be found in Appendix F) and Lee et al. (2021) show that GeometricHedge.P (Bartlett et al., 2008) with John’s exploration meets the condition for linear bandits with $g(\delta)=\mathcal{O}\big{(}d\log(|{\mathcal{A}}|/\delta)\big{)}$. Hence, feeding those algorithms to Algorithm 2 yields the following:
Theorem 3.5.
For any fixed $\delta\in(0,1)$, if Algorithm 2 uses
[…]
Theorem 3.5 suggests that Algorithm 2 enables Exp3.P and GeometricHedge.P to achieve near-optimal ULI guarantees for MAB and linear bandits, respectively. Moreover, their ULI guarantees are as good as those of conventional elimination-based algorithms (refer to Theorem 3.2).
4 Achieving ULI for Linear Bandits in Large Arm-Space
@@ 4.1 Main Algorithm and Main Results
The starting point of our algorithm design is the phased elimination (PE) algorithm (Lattimore et al., 2020, Algorithm 12, Chapter 22). However, PE is in general not feasible when the arm space is large (e.g., a continuous space). In this section, we present a carefully-designed algorithm to address the new challenges from large arm-spaces.
Issues of PE for large arm-space. PE needs to (i) compute (approximate) $G$-optimal design whose complexity scales linearly with $|{\mathcal{A}}|$ and (ii) compare the empirical mean of each arm, both of which are impossible when the arm set is infinite, e.g., ${\mathcal{A}}$ is a ball. A natural idea is to discretize ${\mathcal{A}}$, e.g., constructing a $\epsilon$-net, but the computational complexity has an exponential dependence on $d$, and the optimal arm does not necessarily lie in the net, which prevents the convergence to the optimal arm, i.e., $\lim_{t\to+\infty}\Delta_{A_{t}}=0$.
@@
Before describing the algorithm, we introduce the definition of barycentric spanner (Awerbuch and Kleinberg, 2008).
@@
Algorithm 3 PE with adaptive barycentric spanner
Initialize: $\widehat{\theta}_{1}=\vec{0}\in\mathbb{R}^{d}$, $T_{0}=1$ and ${\mathcal{B}}_{0}=\{e_{1},\ldots,e_{d}\}$.
for $m=1,2,\ldots$ do
@@
Then, the algorithm assigns $\pi_{m}(a)=\frac{1}{d}$ for each $a\in{\mathcal{B}}_{m}$. In fact, if we add $\cup_{i\in{\mathcal{I}}_{m}}\frac{e_{i}}{\sqrt{T_{m}}}$ (see line 4 of Algorithm 4) back to ${\mathcal{B}}_{m}$ and denote the new set by ${\mathcal{B}}_{m}^{\prime}$, then $\pi_{m}:{\mathcal{B}}_{m}^{\prime}\to\frac{1}{d}$ forms an optimal design. It is noteworthy that our algorithm only plays arms in ${\mathcal{B}}_{m}$ because these added elements do not necessarily exist in the arm set ${\mathcal{A}}_{m}$. As each added element $\frac{e_{i}}{\sqrt{T_{m}}}$ is close to zero, even if we only play arms in ${\mathcal{B}}_{m}$, we can still acquire accurate estimations uniformly over ${\mathcal{A}}_{m}$ (refer to Appendix G.3 for details):
$\forall a\in{\mathcal{A}}_{m},\quad\left\|{a}\right\|_{V_{m}^{-1}}\leq C\cdot d/\sqrt{T_{m}},$ (6)
@@
For any fixed $\delta\in(0,1)$, Algorithm 3 achieves the ULI guarantee with function
[…]
Moreover, in each phase, the number of calls to the optimization oracle given in Definition 4.4 is
$\mathcal{O}\left(d^{3}\log_{C}d\right).$

### v2

@@ 2.3 New Metric: Uniform Last-Iterate Guarantee
To address the aforementioned issue, we introduce a new metric, formally defined below.
@@
An algorithm is ULI for some $\delta\in(0,1)$ if there exists a positive-valued function $F_{\text{ULI}}(\cdot,\cdot)$, such that
$\mathbb{P}\left(\forall t\in\mathbb{N}:\Delta_{t}\leq F_{\text{ULI}}(\delta,t)\right)\geq 1-\delta,$
where $F_{\text{ULI}}(\delta,t)$ is polynomial in $\log(1/\delta)$ and proportional to the product of power functions of $\log t$ and $1/t$ (e.g., $F_{\text{ULI}}(\delta,t)=\mathtt{polylog}(1/\delta)(\log t)^{\kappa_{1}}t^{-\kappa_{2}}$ for some $\kappa_{1},\kappa_{2}\geq 0$).
According to Definition 2.5, the instantaneous suboptimality of any algorithm with the ULI guarantee can be bounded by a function $F_{\text{ULI}}(\delta,t)$. Moreover, $F_{\text{ULI}}(\delta,t)$ decreases monotonically for large $t$ if its power on $1/t$ is strictly positive, which captures the convergence rate of $\Delta_{t}$.
Note that the convergence rate of $\Delta_{t}$ is mainly determined by the power on $1/t$ in $F_{\text{ULI}}$. Moreover, as we will show shortly, an algorithm with the ULI guarantee automatically has a small regret bound. We therefore have the following lower bound on $F_{\text{ULI}}$.
@@
For any bandit algorithm that achieves ULI guarantee for some $\delta$ with function $F_{\text{ULI}}(\delta,t)$, there exists a MAB instance such that $F_{\text{ULI}}(\delta,t)=\Omega\left(t^{-1/2}\right)$.
@@
Suppose an algorithm achieves the ULI guarantee for some $\delta$ with function $F_{\text{ULI}}(\delta,t)=\mathtt{polylog}(t/\delta)\cdot t^{-\kappa}$ where $\kappa\in(0,1)$. Then, we have,
  • the algorithm is uniform-PAC with function $F_{\text{UPAC}}(\delta,\epsilon)=\mathcal{O}\big{(}\epsilon^{-\frac{1}{\kappa}}\cdot\mathtt{polylog}(\delta^{-1}\epsilon^{-1})\big{)}.$
@@
    $\mathcal{O}\left(\min\left\{\mathtt{polylog}(T/\delta)\cdot T^{1-\kappa},\Delta^{1-1/\kappa}\mathtt{polylog}^{2}\left((\delta\Delta)^{-1}\right)\right\}\right),$
@@
According to Theorem 2.7, if an algorithm is with near-optimal ULI guarantee (i.e., $\kappa=\frac{1}{2}$), then it implies the near-optimality for uniform-PAC bound (the first bullet point) and anytime sublinear high-probability regret bound (the second bullet point). On the other hand, an algorithm with near-optimal uniform-PAC bound does not necessarily enjoy a near-optimal ULI guarantee as shown in Section 3.2. The proof of Theorem 2.7 can be found in Appendix B.
Remark 2.8.
Although a near-optimal ULI guarantee implies an anytime sublinear high-probability regret bound, it cannot give an anytime sublinear expected regret bound. This is because any algorithm with ULI guarantee is also uniform-PAC, but (Dann et al., 2017, Theorem 1) implies that no algorithm can be uniform-PAC and achieve anytime sublinear expected regret bound simultaneously.
3 Achieving Near-Optimal ULI in Bandits with Finite Arm-Space
In this section, we answer the question whether ULI is achievable for bandit problems. To this end, we examine three common types of bandit algorithms, including elimination-based algorithms, optimistic algorithms, and high-probability adversarial algorithms, in the finite arm setting, i.e., $|{\mathcal{A}}|=K$.
@@ 3.1 Elimination Framework Achieving Near-Optimal ULI Guarantee
To examine whether the elimination-type algorithms can achieve the ULI guarantee, we first provide an elimination framework in Algorithm 1 that ensures the ULI guarantee, and we then show that most elimination-based algorithms fall into this framework. The following result shows that with a proper function $f$ and a positive constant $\beta$, such an elimination framework ensures the ULI guarantee.
@@
For any given $\delta\in(0,1)$, if there exists function $f(\delta,t)=t^{-\kappa}\mathtt{polylog}(t/\delta)$ for some $\kappa\in(0,1)$ and $\exists\beta>0$, such that with probability $1-\delta$, Eq. (1) holds for all $t$, then algorithm is ULI with $F_{\text{ULI}}(\delta,t)=\mathcal{O}\left(f(\delta,t)\right)$.
@@
       Select an active set ${\mathcal{A}}_{t}$ based on available observations as ($a^{*}$ is one of optimal arms)
${\mathcal{A}}_{t}\subseteq\left\{a\in{\mathcal{A}}_{t-1}:\Delta_{a}\leq\beta\cdot f\left(\delta,t\right)\right\}\cup\left\{a^{\star}\right\}.$ (1)
Play an arm $A_{t}\in{\mathcal{A}}_{t}$ and observe reward $X_{t,A_{t}}$.
Theorem 3.1 suggests that Eq. (1) is a sufficient condition for elimination-based algorithms to achieve the ULI guarantee. Now, we show that existing elimination algorithms indeed fall into such a framework. We here consider successive elimination (SE) and phased elimination (PE). Notice that we consider SE only for the MAB setting (called SE-MAB, e.g., Algorithm 3 in (Even-Dar et al., 2006)) but PE for both the MAB setting (called PE-MAB, e.g., exercise 6.8 in (Lattimore and Szepesvári, 2020)) and the linear bandit setting (called PE-L, e.g., Algorithm 12 of Chapter 22 in (Lattimore and Szepesvári, 2020)). Since those algorithms are standard, we defer their pseudocodes to Appendix D. Given Theorem 3.1, the following results show that the elimination framework can be instantiated by these algorithms with proper functions and therefore they achieve the ULI.
@@
For any fixed $\delta\in(0,1)$, elimination framework in Algorithm 1 can be instantiated by
  • SE-MAB for MAB to achieve the ULI with $F_{\text{ULI}}(\delta,t)=\mathcal{O}\big{(}t^{-\frac{1}{2}}\sqrt{K\log(\delta^{-1}Kt)}\big{)}$.
  • PE-MAB for MAB to achieve the ULI with $F_{\text{ULI}}(\delta,t)=\mathcal{O}\big{(}t^{-\frac{1}{2}}\sqrt{K\log(\delta^{-1}K\log(t+1))}\big{)}$.
  • PE-L for linear bandits to achieve ULI with $F_{\text{ULI}}(\delta,t)=\mathcal{O}\big{(}t^{-\frac{1}{2}}\sqrt{d\log(\delta^{-1}K\log(t+1)\log d)}\big{)}$.
Achieving ULI by adversarial bandit algorithms. Traditional elimination-based algorithms, including the ones mentioned above, typically require a carefully designed exploration strategy which is non-trivial even for linear bandits. Here, we provide an alternative way to achieve ULI by employing adversarial bandit algorithms to explore and then conduct the elimination. As shown in Appendix F, all adversarial bandit algorithms for both MAB and linear bandits that meet a certain condition can naturally be used to achieve the ULI guarantees similar to those of traditional elimination-based algorithms.
@@ 3.2 Lower Bound for Optimistic Algorithms
In this section, we present a lower bound to show that optimistic algorithms cannot achieve near-optimal ULI guarantee. The procedure of optimistic algorithms is summarized as follows. After playing each arm once, at each round $t$, the algorithm plays an arm $A_{t}$ that satisfies
@@
where $N_{a}(t)$ is the number of plays of arm $a$ before round $t$, $\widehat{\mu}_{a}(N_{a}(t))=\frac{\sum_{s=1}^{t-1}X_{s,A_{s}}\mathbb{I}{\left\{{A_{s}=a}\right\}}}{N_{a}(t)}$ is the empirical mean of arm $a$ after $N_{a}(t)$ times play, and $U_{\delta}(N_{a}(t))$ is a positive bonus function which encourages the exploration.
We first consider optimistic algorithms e.g., upper confidence bound (UCB) (Lattimore and Szepesvári, 2020, Algorithm 3, Chapter 7), which enjoy (near)-optimal regret bounds. This type of algorithms typically uses the bonus function in the form of $\sqrt{\log(2t^{2}/\delta)/N_{a}(t)}$555 We slightly adjust the bonus function to ensure that with probability at least $1-\delta$, the confidence bound holds for all $t\in\mathbb{N}$.. However, the $\log t$ term forces the algorithm to play suboptimal arms infinitely often, and thus they cannot achieve the ULI guarantee (refer to Appendix E.3 for a detailed proof). Similarly, other variants (Audibert and Bubeck, 2010; Degenne and Perchet, 2016) with $\log t$ term in bonus function, also cannot achieve the ULI guarantee.
We then consider another optimistic-type algorithm, lil’UCB (Jamieson et al., 2014) which obtains the order-optimal instance-dependent sample complexity and avoids $\log(t)$ term in $U_{\delta}(N_{a}(t))$. The bonus function of lil’UCB is as $\sqrt{\log\left(\delta^{-1}\log_{+}\left(N_{a}(t)\right)\right)/N_{a}(t)}$ where $\log_{+}(x)=\log\left(\max\left\{x,e\right\}\right)$. Our main result for lil’UCB is presented as follows. The full analysis of lil’UCB is deferred to Appendix E.
@@
There exists a constant $\alpha\in(0,1)$ that for all $\Delta\in(0,\alpha)$, running lil’UCB on the two-armed bandit instance with deterministic rewards and arm gap $\Delta$ gives $\exists t=\Omega\left(\Delta^{-2}\right)$ such that
$\Delta_{t}=\Omega\left(t^{-\frac{1}{4}}\sqrt{\log\log\left(\Delta^{-2}\log(\delta^{-1})\right)+\log(\delta^{-1})}\right).$
Theorem 3.3 shows that lil’UCB is not near-optimal ULI, but it is unclear whether it can achieve the ULI guarantee. Recall from Theorem 3.2 that the elimination-based algorithms ensure that with high probability, $\forall t$, $\Delta_{t}=\widetilde{\mathcal{O}}\left(t^{-\nicefrac{{1}}{{2}}}\right)$. Theorem 3.3 suggests that the convergence rate of lil’UCB is strictly worse than that of elimination-based algorithms, even if it enjoys a near-optimal cumulative performance (Jamieson et al., 2014). In fact, this lower bound holds for all optimistic algorithms as long as the bonus function is in a similar form.
Remark 3.4 (ULI is strictly stronger than uniform-PAC).
For lil’UCB, the number of times playing suboptimal arms is finite with an order-optimal instance-dependent sample complexity, which implies that lil’UCB is near-optimal uniform-PAC. Therefore, Theorem 3.3 also shows that an algorithm with near-optimal uniform-PAC does not necessarily enjoy near-optimal ULI guarantee.
4 Achieving Near-Optimal ULI for Linear Bandits in Large Arm-Space
@@ 4.1 Main Algorithm and Main Results
The starting point of our algorithm design is the phased elimination (PE) algorithm (Lattimore et al., 2020, Algorithm 12, Chapter 22). However, PE in general is not feasible when the arm space is large (e.g., a continuous space). In this section, we present a carefully-designed algorithm to address the new challenges from large arm-spaces.
Algorithm 2 PE with adaptive barycentric spanner
Initialize: $\widehat{\theta}_{1}=\{0,\ldots,0\}\in\mathbb{R}^{d}$, $T_{0}=1$ and ${\mathcal{B}}_{0}=\{e_{1},\ldots,e_{d}\}$.
for $m=1,2,\ldots$ do
      65432 Set $T_{m}=256C^{4}\cdot\frac{d^{3}}{4^{-m}}\log\left(\delta^{-1}d^{3}4^{m}\right)$. Invoke Algorithm 12 with $({\mathcal{A}},m,{\mathcal{B}}_{m-1},T_{m},C,\widehat{\theta}_{m})$ to find a $C$-approximate barycentric spanner ${\mathcal{B}}_{m}$ for active arm set ${\mathcal{A}}_{m}$ where ${\mathcal{A}}_{m}$ is in Eq. (3). Set $\pi_{m}(a)=\frac{1}{d}$ for each $a\in{\mathcal{B}}_{m}$. Play each arm $a\in{\mathcal{B}}_{m}$ for $n_{m}(a)=\left\lceil T_{m}\pi_{m}(a)\right\rceil$ times. Compute $V_{m}=I+\sum_{a\in{\mathcal{B}}_{m}}n_{m}(a)aa^{\top}$ and $\widehat{\theta}_{m+1}=V_{m}^{-1}\sum_{t\in{\mathcal{T}}_{m}}A_{t}X_{t,A_{t}}$ where ${\mathcal{T}}_{m}$ is a set that contains all rounds in phase $m$.
Issues of PE for large arm-space. PE needs to (i) compute (approximately) $G$-optimal design whose complexity scales linearly with $|{\mathcal{A}}|$ and (ii) compare the empirical mean of each arm, both of which are impossible when the arm set is infinite, e.g., ${\mathcal{A}}$ is a ball. A natural idea is to discretize ${\mathcal{A}}$, e.g., constructing a $\epsilon$-net, but the computational complexity has an exponential dependence on $d$, and the optimal arm does not necessarily lie in the net, which prevents the convergence to the optimal arm.
@@
The definition of barycentric spanner (Awerbuch and Kleinberg, 2008) is presented as follows.
@@
Why adaptive barycentric spanner? The primary reason that the non-adaptive barycentric spanner technique (Awerbuch and Kleinberg, 2008) fails to work for PE is that it requires the knowledge of the space that the active arm set spans. However, acquiring such knowledge is often difficult because the active arm set, which may span a proper linear subspace of $\mathbb{R}^{d}$, varies for different phases and the number of active arms could be infinite. Our algorithm shown in Algorithm 12 (whose pseudocode is deferred to Appendix G.2 due to space limit) can identify a barycentric spanner for active arm set adaptively for each phase, even if they do not span $\mathbb{R}^{d}$.
@@
Then, the algorithm assigns $\pi_{m}(a)=\frac{1}{d}$ for each $a\in{\mathcal{B}}_{m}$. In fact, if we add $\cup_{i\in{\mathcal{I}}_{m}}\frac{e_{i}}{\sqrt{T_{m}}}$ (refer Appendix G for $e_{i}$ and ${\mathcal{I}}_{m}$) back to ${\mathcal{B}}_{m}$ and denote the new set by ${\mathcal{B}}_{m}^{\prime}$, then $\pi_{m}:{\mathcal{B}}_{m}^{\prime}\to\frac{1}{d}$ forms an optimal design. It is noteworthy that our algorithm only plays arms in ${\mathcal{B}}_{m}$ as these added elements do not necessarily exist in the arm set ${\mathcal{A}}_{m}$. As each added element $\frac{e_{i}}{\sqrt{T_{m}}}$ is close to zero, even if we only play arms in ${\mathcal{B}}_{m}$, we can still acquire accurate estimations uniformly over ${\mathcal{A}}_{m}$ (refer to Appendix G.6 for details):
$\forall a\in{\mathcal{A}}_{m}:\quad\left\|{a}\right\|_{V_{m}^{-1}}=\sqrt{a^{\top}V_{m}^{-1}a}\leq C\cdot d/\sqrt{T_{m}},$ (4)
@@
For any fixed $\delta\in(0,1)$, Algorithm 2 achieves the ULI guarantee with function $F_{\text{ULI}}(\delta,t)=\mathcal{O}\big{(}t^{-\frac{1}{2}}\sqrt{d^{3}\log(\delta^{-1}dt)}\big{)}$. Moreover, in each phase, the number of calls to the optimization oracle given in Definition G.2 is $\mathcal{O}\left(d^{3}\log_{C}d\right)$.